OpenStudy (okdutchman7):

Please help. WILL MEDAL! How many distinguishable permutations can formed from the letters of the word sassafras? @pooja195 @Mehek14 @nincompoop @jigglypuff314 @KendrickLamar2014 @satellite73 @zepdrix @rebeccaxhawaii @bahrom7893 @Elsa213 @xapproachesinfinity @dmezzullo @Owlcoffee

OpenStudy (okdutchman7):

I really don't know where to begin.

OpenStudy (okdutchman7):

s: 4 a: 3 f: 1 r: 1

OpenStudy (loser66):

We have formula for this, right?

OpenStudy (okdutchman7):

I don't know what it is.

OpenStudy (loser66):

\(\dfrac{total!}{s!a!f!r!}\)

OpenStudy (okdutchman7):

\[\frac{ 9! }{ 4! 3! }\]

OpenStudy (loser66):

yup, but you have to put 1!1! for f and r to show that they are there.

OpenStudy (okdutchman7):

\[\frac{ 362,880 }{ 24\times6 }\]

OpenStudy (okdutchman7):

but they are just 1 correct?

OpenStudy (loser66):

I didn't calculate it. calculation is the easiest step. I think you can handle it. Yup for the last question.

OpenStudy (okdutchman7):

\[\frac{ 362,880 }{ 144}\]

OpenStudy (okdutchman7):

2520 permutations

OpenStudy (okdutchman7):

Thanks!

OpenStudy (loser66):

ok